OpenStudy (anonymous):

LET f(x) = x^4 + ax^2 + b . The graph of f has a relative maximum at (0,1) and and inflection point at x=1 . Find the values of a and b

OpenStudy (danjs):

In order to do that, we should calculate the derivatives of f(x) first

OpenStudy (danjs):

do you remember the power rule?

OpenStudy (anonymous):

4x^3 + 2ax ?

OpenStudy (danjs):

yep

OpenStudy (danjs):

At a maximum value, the slope of a tangent will be zero, and the derivative is the slope of the tangent, set f '(x) = 0

OpenStudy (anonymous):

k hold on

OpenStudy (danjs):

\[f ' (x) = 0 = 4x^2+ 2ax\] now calculate the second derivative

OpenStudy (anonymous):

I set it to 0 and I am solving for what because I have a and x

OpenStudy (anonymous):

8x + 2a

OpenStudy (danjs):

right, \[f ''(x) = 8x + 2a \] f ''(x) = 0 , when x = 1

OpenStudy (danjs):

the inflection point is where the second derivative = 0, the graph switches concavity

OpenStudy (anonymous):

yeah so how do I find that , I make the second derrivative equal to 0?

OpenStudy (danjs):

exactly, f ''(1) = 0

OpenStudy (anonymous):

so I plug 1 in for a ?

OpenStudy (anonymous):

or x

OpenStudy (danjs):

no, the inflection point occurs when x = 1

OpenStudy (danjs):

we are trying to find a

OpenStudy (anonymous):

but I have to find the values of a and b

OpenStudy (danjs):

yes, we use both equations, the first derivative, and the second, 2 equations, 2 unknowns a and b

OpenStudy (anonymous):

ok so we found the derrivatives but what is the value of a and b?

OpenStudy (danjs):

actually, to get b, you can just use the regular function f(x) at point (0,1)

OpenStudy (danjs):

f(0) = 1 = 0 + a*0 + b b = 1

OpenStudy (anonymous):

ok that was the normal equation but what about a ? do I use the first or second derivative

OpenStudy (danjs):

either one will be fine, since you are only finding one variable, i would go with the second derivative

OpenStudy (anonymous):

look at all the ghosts

OpenStudy (danjs):

BOO!!!!

OpenStudy (danjs):

it must be North Korea

OpenStudy (anonymous):

oh now they are going away bye bye ghosts

OpenStudy (anonymous):

what are ghosts ? and I do not know how to get a

OpenStudy (danjs):

f '' (x) = 8x + 2a at x = 1 inflection f''(x) = 0 f ''(1) = 0 = 8(1) + 2a

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

oh

OpenStudy (danjs):

wait, we messed up the second derivative...the first function is to the 4th power

OpenStudy (anonymous):

4x^3

OpenStudy (danjs):

f(x) = x^4 + ax^2 + b f ' (x) = 4x^3 + 2ax f '' (x) = 12x^2 + 2a

OpenStudy (danjs):

f ''(x) = 0 at x=1 inflection point f ''(1) = 12(1)^2 + 2a = 0 12 + 2a = 0

OpenStudy (danjs):

a = -6

OpenStudy (anonymous):

thank you so much!

OpenStudy (danjs):

you understand what they want you to know? f ' (x) = 0 ...maximum/minimum values f '' (x) = 0 ...inflection points

OpenStudy (anonymous):

yes

OpenStudy (danjs):

but from the function in this problem, when x=0 from the point (0,1) , you can find b from just f(x) by letting x=0

OpenStudy (danjs):

x = 0, f(0) = b

OpenStudy (danjs):

some probs, you might have to use the first derivative as a maximum or minimum to find thevariable, if the function doesnt simplify like that

OpenStudy (anonymous):

alright , I got it (: thank you !!

OpenStudy (danjs):

anytime, pretty fun prob there

OpenStudy (danjs):

i would graph it, and use a slider for a and b, you can change a and b on the fly and see what the graph does

OpenStudy (danjs):

https://www.desmos.com/calculator

OpenStudy (danjs):

https://www.desmos.com/calculator